OpenStudy (anonymous):

Factor Completely: Sin^4 - Cos ^4

OpenStudy (anonymous):

Explain please

OpenStudy (anonymous):

\[a^4-b^4=(a^2+b^2)(a+b)(a-b)\]

OpenStudy (anonymous):

in this case the first term drops out because \[\sin^2(x)+\cos^2(x)=1\]

OpenStudy (anonymous):

what is the first term?

OpenStudy (anonymous):

the a^2+b^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why does it drop out? i dont really understand

OpenStudy (anonymous):

replace \(a=\sin(x)\) and \(b=\cos(x)\) to get your answer

OpenStudy (anonymous):

do you know what \(\sin^2(x)+\cos^2(x)\) is?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

because of a property

OpenStudy (anonymous):

that is why it drops out. the first term is 1, so you don't need to write it

OpenStudy (anonymous):

ohhhh ok i didnt know what you meant by drop out but now i see

OpenStudy (anonymous):

so then it is ( sin^2(x)+cos^2(x))* (sin^2(x)-cos^2(x))

OpenStudy (anonymous):

IS THE ANSWER 2 or 1?

OpenStudy (mertsj):

\[\sin ^4x-\cos ^4x=(\sin ^2x+\cos ^2x)( \sin ^2x-\cos ^2x)=1(\sin x+\cos x)(\sin x-\cos x)\]

OpenStudy (mertsj):

The answer is (sinx+cosx)(sinx-cosx)

OpenStudy (anonymous):

but isnt that sin^2+cos^2

OpenStudy (anonymous):

when you foil it

OpenStudy (anonymous):

so then 1*1=1...?

OpenStudy (callisto):

I'm thinking this... \[sin^4x - cos ^4x\]\[=(sin^2x - cos ^2x)(sin^2x+cos^2x) \]\[=(sin^2x - cos ^2x) (1)\]\[=-(cos ^2x-sin^2x)\]\[=-cos2x\]

OpenStudy (mertsj):

is cosx(-cosx)=+cos^2x?????

OpenStudy (anonymous):

oh.. no...

OpenStudy (anonymous):

callisto, im confused how does cos^2x-sin^2=-cos^2

OpenStudy (callisto):

Not ^2 ... It's 2... \(cos2x = cos^2x-sin^2x\) It's an identity...

OpenStudy (anonymous):

? I don't think i know that identity. Not in my book either..

OpenStudy (anonymous):

but so far, ihave 1*sin^2-cos^2...

OpenStudy (callisto):

then @Mertsj 's answer is what you want ... but there is such an identity!!!! use difference of two squares again... a^2 - b^2 = (a+b)(a-b) a= sinx , b= cosx

OpenStudy (anonymous):

so the answer is the foiled version of Mertsj? sin^2-cos^2?

OpenStudy (callisto):

shouldn't you factor it?

OpenStudy (anonymous):

(sinx+cosx)(sinx-cosx) is what he had

OpenStudy (anonymous):

so foil? do you mean the same thing?

OpenStudy (callisto):

this is what you should have :)

OpenStudy (anonymous):

but that is the same as

OpenStudy (anonymous):

Sin^2-sincos+sincos-cos^2. Sincos cancels, leaving sin^2-cos^2

OpenStudy (anonymous):

so shouldnt it be sin^2-cos^2?

OpenStudy (callisto):

Yes, they are the same, sin^2x- cos^2x = (sinx+cosx)(sinx-cosx) But!!!! (sinx+cosx)(sinx-cosx) is the factorised form, this is what required in the question... while sin^2x - cos^2x is not

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

can you help me with a quick one? i don't understand how this one is done...

OpenStudy (callisto):

Hmm... can I eat my orange first? I'm not a good multitasker...

OpenStudy (anonymous):

yes of course hahaha

OpenStudy (anonymous):

i will help you practice multitask :)